### v1

# HumanSplat: Generalizable Single-Image Human Gaussian Splatting with Structure Priors

Abstract
Despite recent advancements in high-fidelity human reconstruction techniques, the requirements for densely captured images or time-consuming per-instance optimization significantly hinder their applications in broader scenarios. To tackle these issues, we present HumanSplat that predicts the 3D Gaussian Splatting properties of any human from a single input image in a generalizable manner. In particular, HumanSplat comprises a 2D multi-view diffusion model and a latent reconstruction transformer with human structure priors that adeptly integrate geometric priors and semantic features within a unified framework. A hierarchical loss that incorporates human semantic information is further designed to achieve high-fidelity texture modeling and better constrain the estimated multiple views. Comprehensive experiments on standard benchmarks and in-the-wild images demonstrate that HumanSplat surpasses existing state-of-the-art methods in achieving photorealistic novel-view synthesis. Project page: https://humansplat.github.io/
1 Introduction
Realistic 3D human reconstruction is a fundamental task in computer vision with widespread applications in numerous fields, including social media, gaming, e-commerce, telepresence, etc. Previous works for single-image human reconstruction can be broadly categorized into explicit and implicit approaches. Explicit methods pymaf2021 ; PIXIE:2021 ; alldieckTex2ShapeDetailedFull2019 ; alldieck2018dv , such as those based on parametric body models like SMPL SMPL:2015 ; SMPL-X:2019 , estimate human meshes by directly optimizing parameters and clothing offset to fit the observed image. However, these methods often struggle with complex clothing styles and require lengthy optimization. Implicit methods saito2019pifu ; saito2020pifuhd ; alldieck2022photorealistic ; hu2023sherf ; huang2022elicit represent humans using continuous functions, such as occupancy saito2019pifu , SDF park2019deepsdf , and NeRF mildenhall2021nerf . They are good at modeling flexible topology but are limited in scalability and efficiency due to the high computational cost of training and inference. Recent advances in 3D Gaussian Splatting kerbl20233d (3DGS) have provided a balance between efficiency and rendering quality for reconstructing detailed 3D human models, which however rely on multi-view zielonka2023drivable ; moreau2023human ; qian20233dgs or monocular video li2024gaussianbody ; hu2023gauhuman ; hu2023gaussianavatar inputs. Recent popular human reconstruction studies TECH:2023 ; HumanSGD:2023 ; humannorm:2023 ; HAVEFUN:2023 focus on the score distillation sampling (SDS) technique SDS:2022 to lift 2D diffusion priors to 3D, but time-consuming optimization (e.g., 2 hours) is required for each instance. Some generalizable and large-reconstruction-model-based works hu2023sherf ; ActorsNeRF:2023 ; HumanLRM:2024 ; GPS:2023 can directly generalize the regression of 3D representations but either disregard human priors or require multi-view inputs, limiting the stability and feasibility in downstream applications.
In this paper, we propose a novel generalizable approach HumanSplat for single-image human reconstruction by introducing a generalizable Gaussian Splatting framework with a pre-trained 2D video diffusion model and well-designed 3D human structure priors. Different from the existing human 3DGS methods, our approach directly infers Gaussian properties from a single image, obviating the necessity for per-instance optimization or densely captured images and empowering to generalize effectively across diverse scenes while efficiently delivering high-quality reconstructions.
The key insight behind our method is to reconstruct Gaussian properties from the diffusion latent space in a generalizable end-to-end architecture, integrating a 2D generative diffusion model as appearance prior and a human parametric model as structure prior. Specifically, facing this under-constraint reconstruction problem from single-view input with extensive invisible regions, we first leverage a 2D video diffusion model (denoted as novel-view synthesizer) to hallucinate the unseen parts of clothed humans. Then a generalizable latent reconstruction transformer is proposed to enable interaction among the generated diffusion latent embeddings and the geometric information of the structured human model, enhancing the quality of 3DGS reconstruction. During interactions using inaccurate human prior like the SMPL model, a projection strategy is designed to balance the robustness and flexibility, where we compensate for inaccuracies by projecting 3D tokens into 2D space and conducting searches within adjacent windows using project-aware attention. Another specific challenge of human reconstruction is to capture fine detail in visually sensitive areas, such as the face and hand. To address this problem, we exploit the semantic cues from the structure priors and propose semantics-guided objectives to further promote the fine details reconstruction quality. By amalgamating these crucial components into a unified framework, our method achieves state-of-the-art performance in striking the right balance between quality and efficiency.
In summary, the main contributions of our work are as follows:
  • We propose a novel generalizable human Gaussian Splatting network for high-fidelity human reconstruction from a single image. To the best of our knowledge, it is the first to leverage latent Gaussian reconstruction with a 2D generative diffusion model in an end-to-end framework for efficient and accurate single-image human reconstruction.
  • We integrate structure and appearance cues within a universal transformer framework by leveraging human geometry priors from the SMPL model and human appearance priors from the 2D generative diffusion model. Geometric priors stabilize the generation of high-quality human geometry, while appearance prior helps hallucinate unseen parts of clothed humans.
  • We enhance the fidelity of reconstructed human models by introducing semantic cues, hierarchical supervision, and tailored loss functions. Extensive experiments demonstrate that our method achieves state-of-the-art performance, surpassing existing methods.
Figure 1: Our method achieves state-of-the-art rendering quality while maintaining the fastest runtime. (a) Qualitative results: LGM LGM:2024 and GTA GTA:2024 are generalizable but in lower quality, TeCH TECH:2023 exhibits issues with multi-face rendering and is time-consuming. In contrast, our method achieves higher fidelity in a much shorter time. (b) Performance and runtime comparison: metrics are evaluated on the challenging Twindom dataset.
2 Related Work
Single-Image Human Reconstruction. Single-image human reconstruction methods fall into explicit and implicit approaches. Explicit methods use parametric body models SMPL:2015 ; SMPL-X:2019 to estimate minimally clothed human meshes pymaf2021 ; PIXIE:2021 ; Kanazawa2018_hmr ; pare ; pymafx2023 ; li2023niki ; chen2022learning ; VIBE ; Kolotouros2019_spin , adding clothing via 3D offsets alldieckTex2ShapeDetailedFull2019 ; alldieck2018dv ; alldieck2018videoavatar ; alldieck2019peopleInClothing ; zhu2019hierarchMeshDeform ; lazova2019textures360 or garment templates bhatnagar2019multiGarmentNet ; jiang2020bcnet ; feng2023learning . However, these methods are constrained by topology, particularly with loose clothing. Implicit methods saito2019pifu ; saito2020pifuhd ; alldieck2022photorealistic ; han2023high ; albahar2023humansgd ; hu2023sherf ; huang2022elicit ; HumanLRM:2024 ; Zhang_2024_CVPR , utilizing representations like occupancy, signed distance fields (SDF) and Neural Radiance Fields (NeRF), offer flexibility with topology, allowing for accurate depiction of 3D clothed humans. Recent methods he2020geoPifu ; corona2023s3f ; xiu2022icon ; zheng2021pamir ; huangARCHAnimatableReconstruction2020 ; heARCHAnimationReadyClothed2021 ; cao2022jiff ; liao2023car ; yang2023dif ; cao2023sesdf ; zhang2023globalcorrelated ; robustfusion ; su2022robustfusionPlus ; xiu2022econ ; Jiang_2023_CVPR ; zheng2024ohta combine implicit representation with explicit human priors for better robustness. Among these, GTA zhang2023globalcorrelated uses transformers with fixed embeddings for translating image features into 3D tri-plane features. TeCH huang2024tech employs diffusion-based models for visualizing unseen areas but needs extensive optimization and accurate SMPL-X models. Another approach, HumanSGD albahar2023humansgd , uses diffusion models for texture inpainting but struggles with mesh inaccuracies from saito2020pifuhd . NeRF-based methods like SHERF hu2023sherf generate human NeRFs from single images using hierarchical features, while ELICIT huang2022elicit leverages CLIP CLIP:2021 for contextual understanding. These NeRF-based methods produce high-quality images but often struggle with detailed 3D mesh generation and require extensive optimization time.
Human Gaussian Splatting. While implicit representations like SDF or NeRF struggle with balance optimization efficiency and rendering quality, the high efficiency of 3DGS kerbl3Dgaussians has advanced 3D human creation. Recent 3DGS methods have targeted multi-view videos, monocular video sequences, and sparse-view input. Multi-view video methods like D3GA zielonka2023drivable , HuGS moreau2023human , and 3DGS-Avatar qian20233dgs exploit dense spatial and temporal information for detailed models. Additionally, HiFi4G jiang2023hifi4g combine 3DGS with a dual-graph mechanism to preserve spatial-temporal consistency, ASH pang2023ash utilizes mesh UV parameterization for real-time rendering, and Animatable Gaussians li2023animatable improve garment dynamics via pose projection mechanism and 2D CNNs. The monocular video only provides sufficient observation as humans move, so the monocular human Gaussian Splatting methods li2024gaussianbody ; hu2023gauhuman ; hu2023gaussianavatar ; li2023human101 ; GVA ; SplattingAvatar ; GoMAvatar ; lei2023gart ; HAHA often resort to using LBS for mapping to a canonical space and introducing additional regularization terms. For sparse-view images, GPS-Gaussian GPS:2023 achieves high rendering speeds without per-subject optimization by using a feed-forward way with efficient 2D CNNs encoding from diverse 3D human scan data. In comparison, our work stands out as the first one-shot generalizable human GS method, using only single image input and achieving high-quality reconstruction without any optimization and fine-tuning.
Generalizable Large Reconstrucion Model. Large reconstruction model (LRM) hong2023lrm ; tochilkin2024triposr presents that with sufficient parameters and training datasets, a deterministic feed-forward Transformer vaswani2017attention is capable of learning a triplane feature for volumetric rendering from a single-view image. Targeting general 3D object, TriplaneGaussian zou2023triplane combines 3DGS with triplane as a hybrid 3D representation for fast rendering, where point clouds are first extracted from an image and then projected on triplane features to decode 3DGS attributes. Most recent methods further leverage 2D diffusion models that can generate multi-view images simultaneously shi2023mvdream ; wang2023imagedream ; shi2023zero123++ ; li2023instant3d to provide plausible and abundant inputs for the following reconstruction, so they focus on the sparse multi-view reconstruction task. Instant3D li2023instant3d follows the technique of LRM and produces triplane features from 4 posed images by a Transformer encoder. CRM wang2024crm , MeshLRM wei2024meshlrm and InstantMesh xu2024instantmesh replaces triplane NeRF by FlexiCubes shen2023flexible to extract meshes directly. LGM tang2024lgm , GRM xu2024grm and GS-LRM zhang2024gslrm predict 3DGS attributes from each input pixel szymanowicz2024splatter ; charatan2024pixelsplat and combine the Gaussians from multiple viewpoints as the final 3D output. For human task, CharacterGen peng2024charactergen follows Instant3D li2023instant3d to adopt a two-stage approach: first, generating a 2D multi-view canonical images, then using multi-view SDF reconstruction for cartoon digital characters, without direct interaction between stages. A more compact and elegant method HumanLRM HumanLRM:2024 replaces training data of LRM hong2023lrm from general objects with human data to predict triplane NeRF, showcasing the capacities of the large model. While it relies heavily on the dataset, often leading to issues like missing limbs due to a lack of prior knowledge. In contrast, our 3DGS-based method leverages the latent diffusion and reconstruction model in an end-to-end framework and introduces human priors by addressing their inaccuracy with specific projection strategies, achieving better generalization and stability with even less data.
3 Method
Figure 2: Overview of HumanSplat. (a) Multi-view latent features are first generated by a fine-tuned video diffusion model (Novel View Synthesizer in Sec. 3.3). (b) Then, the Latent Reconstruction Transformer (Sec. 3.4) interacts global latent features (Sec. 3.4.1) and human geometric prior (Sec. 3.4.2). (c) Finally, the semantic-guided objectives (Sec. 3.5) are proposed to reconstruct the final human 3DGS.
3.1 Preliminary
SMPL SMPL:2015 is a widely-used parametric human body model that is created by skinning and blend shapes, and learned from thousands of 3D body scans. The SMPL model SMPL:2015 utilizes shape parameters $\boldsymbol{\beta}\in\mathbb{R}^{10}$ and pose parameters $\boldsymbol{\theta}\in\mathbb{R}^{24\times 3}$ to parameterize the deformation of the human body mesh $\mathcal{M}(\boldsymbol{\beta},\boldsymbol{\theta}):\boldsymbol{\beta}\times\boldsymbol{\theta}\mapsto\mathbb{R}^{6890\times 3}$.
3D Gaussians Splatting kerbl20233d is an recently proposed 3D representation that represents 3D content by a set of $\mathbf{N_{p}}$ colored Gaussians $\mathcal{G}=\{\mathbf{g}_{i}\}_{i=1}^{\mathbf{N_{p}}}$. Each Gaussian $\mathbf{g}_{i}=\exp\left(-\frac{1}{2}(\mathbf{x}-\boldsymbol{\mu}_{i})^{\top}\boldsymbol{\Sigma}_{i}^{-1}(\mathbf{x}-\boldsymbol{\mu}_{i})\right)$ has opacity $\boldsymbol{\sigma}_{i}\in\mathbb{R}$ and color $\mathbf{c}_{i}\in\mathbb{R}^{3}$ attributes, where $\boldsymbol{\mu}_{i}\in\mathbb{R}^{3}$ is its location and $\boldsymbol{\Sigma}_{i}\in\mathbb{R}^{3\times 3}$ implies its scaling $\mathbf{s}\in\mathbb{R}^{3}$ and orientation quaternion $\mathbf{q}\in\mathbb{R}^{4}$ in 3D space. These attributes together $\mathbf{G}\in\mathbb{R}^{\mathbf{N_{p}}\times 14}$ can represent a 3D object, and by projecting to the image plane with opacity-based color composition in the depth order, it can be differentiably rendered into 2D images in real-time.
3.2 Overview
Given a single input image of a human body $\mathbf{I_{0}}\in\mathbb{R}^{H\times W\times 3}$, our task is to reconstruct a 3D representation, i.e., 3DGS in this work, from which novel-view images can be rendered. As illustrated in Fig. 2, our method leverages a 2D diffusion model and a novel latent reconstruction transformer that adeptly integrates 2D appearance priors, human geometric priors, and semantic cues within a universal framework. First, we use SMPL estimator feng2021collaborative ; Pavlakos2019_smplifyx to predict human structure priors $\mathcal{M}$, and CLIP CLIP:2021 to generate the image embedding of the input image $\mathbf{c}$. A latent temporal-spatial diffusion model voleti2024sv3d , referred to as the novel-view synthesizer (Sec. 3.3), is fine-tuned and produces multi-view latent features $\left\{\mathbf{F}_{i}\in\mathbb{R}^{h\times w\times c}\right\}_{i=1}^{\mathbf{N}}$, where $\mathbf{N}$ is the number of views, and $h$, $w$ and $c$ are the height, width and channel of the latent features. Then, we utilize a novel latent reconstruction transformer (Sec. 3.4) that adeptly integrates human geometric prior and latent features within a universal framework, which predicts Gaussian attributes $\mathbf{\mathbf{G}}\coloneqq\left\{(\boldsymbol{\mu}_{i},\boldsymbol{q}_{i},\boldsymbol{s}_{i},\boldsymbol{c}_{i},\boldsymbol{\sigma}_{i})|i=1,...,\mathbf{N_{p}}\right\}$ and rendered into new images, where $\mathbf{N_{p}}$ represents the number of Gaussian points. Furthermore, to achieve high-fidelity texture modeling and better constrain the estimated multiple views, we designed a hierarchical loss that incorporates human semantic priors (See 3.5). Note that during training, the network is trained using 3D scan data to ensure accurate supervision from various viewpoints, where the registered SMPL is fitted via the multi-view version of SMPLify Pavlakos2019_smplifyx . During inference, the model employs PIXIE feng2021collaborative to predict human structure priors $\mathcal{M}$ and synthesize novel views from only the single image input based on the trained model, without any fine-tuning and optimization.
3.3 Video Diffusion Model as Novel-view Synthesizer
We take advantage of the pre-trained video diffusion model SV3D voleti2024sv3d as appearance prior. It utilizes the CLIP CLIP:2021 image embedding $\mathbf{c}$ and latent feature $\mathbf{F}_{0}$ of a input image $\mathbf{I}_{0}$ from a pre-trained VQ-VAE VQGAN:2023 $\mathcal{E}$ as conditions and progressively denoises gaussian noises into temporal-continuous $\mathbf{N}$ multi-view latent features $\{\mathbf{F}_{i}\}_{i=1}^{\mathbf{N}}$ by a spatial-temporal UNet $D_{\theta}$ SVD:2023 with the objective:
$\mathbb{E}_{\epsilon\sim p(\epsilon)}\left[\lambda(\epsilon)\|D_{\theta}(\{\mathbf{F}_{i}^{\epsilon}\}_{i=1}^{\mathbf{N}};\mathbf{c},\mathbf{F}_{0},\epsilon)-\{\mathbf{F}_{i}\}_{i=1}^{\mathbf{N}}\|^{2}_{2}\right],$ (1)
where $\{\mathbf{F}_{i}^{\epsilon}\}$ is the noise-corrupt multi-view latent features, $p(\epsilon)$ is the noise probability distribution and $\lambda(\epsilon)\in\mathbb{R}^{+}$ is the loss weighting term related to the noise level $\epsilon$. Since SV3D is merely pre-trained on the dataset of objects deitke2023objaverse , we further fine-tune it using human datasets to improve its effectiveness in the human reconstruction task. More technical details are available in Appendix A.1.
3.4 Latent Reconstruction Transfomer
Latent reconstruction transformer contains two parts: latent embedding interaction (Sec. 3.4.1) and geometry-aware interaction (Sec. 3.4.2), which seamlessly incorporate latent features from video diffusion and human structure priors into the reconstruction process. The two components together form a cohesive framework for accurately recovering intricate human details.
3.4.1 Latent Embedding Interaction
We utilize a pretrained VQ-VAE VQGAN:2023 $\mathcal{E}$ to encode the input image $\mathbf{I_{0}}$ into a latent representation $\mathbf{F}_{0}=\mathcal{E}(\mathbf{I_{0}})\in\mathbb{R}^{h\times w\times c}$ and combine it with the generated $\left\{\mathbf{F}_{i}\right\}_{i=1}^{\mathbf{N}}$ (Sec. 3.3). Following previous works xu2023dmv3d ; chen2023:ray-conditioning , Plücker ray embeddings plucker1865xvii are leveraged for pose conditioning, where latent representations are concatenated with their corresponding Plücker coordinates along the channel dimension, resulting in a per-pixel pose-conditioned feature map. These latents are then divided into non-overlapping patches dosovitskiy2020image and mapped to $d$-dimensional latent tokens by a linear layer. An intra-attention module is followed to thoroughly extract spatial correlations of latent tokens, as shown in Fig.3. The intra-attention module is implemented using a standard Transformer vaswani2017attention with multiple blocks of multi-head self-attention and feed-forward network (FFN):
$\bar{\mathbf{F}}_{0},\bar{\mathbf{F}}_{1},...,\bar{\mathbf{F}}_{\mathbf{N}}=\text{FFN}(\text{SelfAttention}(\mathbf{F}_{0},\mathbf{F}_{1},...,\mathbf{F}_{\mathbf{N}})).$ (2)
3.4.2 Geometry-aware Interaction
Human Geometric Tokenizer. Given a human geometric model $\mathcal{M}\in\mathbb{R}^{6890\times 3}$, to obtain the feature representation at a position $x\in\mathbb{R}^{3}$, we project $x$ onto input view and computes the feature vector through bilinear interpolation on the feature grids by
$\mathbf{\Pi}_{i}(x)=\mathbf{K}\left(\mathbf{R}\mathcal{M}_{i}+\mathbf{t}\right),$ (3)
where $\mathbf{R,t}$ are camera extrinsic parameters, $\mathbf{K}$ is the intrinsic parameters. After the projection operation, we obtain geometric human prior tokens by concatenating $x$ and $\mathbf{F}_{0}(\mathbf{\Pi}(x))$, and mapping them to $d$ dimension. Similar to Latent Embedding Interaction, an intra-attention module is utilized to interact among these human prior tokens and get geometric features $\bar{\mathbf{H}}_{i}\in\mathbb{R}^{6890\times d}$.
Human Geometry-aware Attention. Empirically, we find that incorporating human priors SMPL:2015 ; SMPL-X:2019 ; SMPLD:2019 (e.g., SMPL model) into human reconstruction faces a significant dilemma of the trade-off between robustness and flexibility GTA:2024 ; SIFU:2024 . On the one hand, the SMPL model provides geometric priors that alleviate common issues (e.g., broken body parts and various artifacts). On the other hand, it struggles to accurately capture and represent a broad spectrum of complex clothing styles, particularly more fluid garments like dresses and skirts. This limitation highlights a fundamental shortcoming in the model’s capacity to accurately depict diverse apparel. As shown in Fig. 3, HumanSplat efficiently utilizes priors while accounting for their inaccuracies by projecting 3D tokens into 2D space and searching in the nearby windows. To tackle the issue of matching redundancy with erroneous human priors, HumanSplat leverages these priors while compensating for their inaccuracies by projecting 3D tokens into 2D space and conducting searches within adjacent windows. Specifically, we introduce project-aware attention within the inter-attention module, utilizing a window $\mathbf{W}(k_{win}\times k_{win})$ and employing masked multi-head cross-attention. Here, the latent features $\left\{\bar{\mathbf{F}}_{i}\right\}$ serve as queries, and the human prior tokens $\left\{\bar{\mathbf{H}}_{i}\right\}$ serve as keys/values. The process can be formulated as follows:
$\tilde{\mathbf{F}}_{i}=\text{FFN}(\text{CrossAttention}_{mask}(\bar{\mathbf{F}}_{i},\bar{\mathbf{H}}_{i})),$ (4)
where feature interactions occur when $\mathbf{\Pi}(\bar{\mathbf{H}}_{i},\mathbf{K},\mathbf{R},\mathbf{t})$ is within the window of $\bar{\mathbf{F}}_{i}$. It is noteworthy that the complexity of the cross-attention operation is reduced from $\mathcal{O}(\mathbf{L}_{F}\times\mathbf{L}_{H})$ to $\mathcal{O}(\mathbf{L}_{F}\times k_{win}^{2})$, compared to the vanilla multi-head Cross-attention setting. Here, $\mathbf{L}_{F}$ and $\mathbf{L}_{H}$ represent the lengths of the latent and human prior tokens respectively.
Figure 3: Illustration of latent reconstruction transformer. It first divides $\mathbf{F}_{0}$ and $\mathbf{F}_{i}$ into non-overlapping patches, which are then processed through an intra-attention module (Sec. 3.4.1). Within the iter-attention module (Sec. 3.4.2), we introduce the project-aware attention with a window $\mathbf{W}(k_{win}\times k_{win})$, and the attributes of 3D Gaussians are decoded with a Conv $1\times 1$ layer.
3.5 Semantics-guided Objectives
From each output token $\tilde{\mathbf{F}}_{i}$, we decode the attributes of pixel-aligned Gaussians $\mathbf{\mathbf{G}}$ in the corresponding patch by a convolutional layer with a $1\times 1$ kernel. These attributes are used to render novel view images through 3D Gaussian Splatting. An ideal training objective should ensure that the rendered outputs closely match the supervised images, striving for overall consistency between the reconstructed renderings and the ground truth. However, human attention tends to be particularly focused on facial regions, where intricate details play a pivotal role in perception. Therefore, in our quest to elevate the fidelity of reconstructed human models, we introduce a paradigm shift in our training objectives.
Hierarchical Loss. Traditional approaches often overlook the semantic richness inherent in human anatomy, leading to reconstructions that lack crucial details and accuracy. To address this, we propose a novel framework that harnesses semantic cues, hierarchical supervision, and tailored loss functions to guide the training process. This ensures that not only are the overall structures of the human body accurately represented, but also that critical areas such as the face are reconstructed with exceptional detail and fidelity. Specifically, by incorporating a human prior model equipped with semantic information, we establish direct correspondences between 3D spatial positions and specific human body parts. Therefore, we can utilize different attention weights for different body parts and render different levels of GT images from different viewpoints to provide hierarchical supervision. This approach significantly facilitates the precise localization of essential body parts, including the head, hands, and arms. The overall loss is defined as a weighted sum of part-specific losses:
$\mathcal{L_{H}}=\frac{1}{l}\frac{1}{n}\sum_{i=1}^{n}\sum_{j=1}^{m}\lambda_{i}\lambda_{j}\mathcal{L}_{Rec}(\mathbf{I}_{\text{part},j},\hat{\mathbf{I}}_{\text{part},j}),$ (5)
where $i\in\left\{1,...,n\right\}$ and $j\in\left\{1,...,m\right\}$ denote different resolution levels and human parts. $\hat{\mathbf{I}}_{\text{part},j}$ and $\mathbf{I}_{\text{part},j}$ are the predicted and ground-truth data for part $j$, $\lambda_{i}$ and $\lambda_{j}$ are weighting factors that reflect the relative importance of each resolution level and each body part, respectively.
Reconstruction Loss. Reconstruction loss $L_{Rec}$ based on the $\mathbf{N}_{\text{render}}$ rendered multi-view images is defined as
$\mathcal{L}_{Rec}=\sum_{i}^{\mathbf{N}_{\text{render}}}\mathcal{L}_{mse}(\hat{\mathbf{I}}_{i},\mathbf{I}_{i})+\lambda_{m}L_{mse}(\hat{\mathbf{M}}_{i},\mathbf{M})+\lambda_{p}\mathcal{L}_{p}(\hat{\mathbf{I}}_{i},\mathbf{I}_{i}),$ (6)
where $\mathbf{I}_{i}$ and $\hat{\mathbf{I}}_{i}$ are the ground-truth images and render images via Gaussian Splatting, $\mathbf{M}_{i}$ and $\hat{\mathbf{M}}_{i}$ are original and rendered foreground mask. $\mathcal{L}_{mse}$ measures the MSE loss and $\mathcal{L}_{p}$ measures the perceptual loss (LIPIS:2018, ). $\lambda_{m}$ and $\lambda_{p}$ are hyperparameters for balancing. The final training objective is a combination of the two loss terms: $\mathcal{L}=\mathcal{L}_{H}+\mathcal{L}_{Rec}$.
4 Experiments
4.1 Implementation Details
[…]
4.2 Comparison
Baselines. We evaluate novel view synthesis using state-of-the-art methods such as GTA GTA:2024 and SIFU SIFU:2024 by rendering textured meshes from multiple angles and calculating the PSNR to ground truth images. We divided 21 scans from Thuman2.0 THuman2.0:2021 and 51 scans from Twindom Twindom for evaluation. To eliminate the bias of the training dataset, we fine-tune the generative model LGM LGM:2024 on the same datasets as ours for a fair comparison. Additionally, we optimized each instance of TeCH TECH:2023 until full convergence for comparison.
[…]
Qualitative Comparison. By incorporating semantics-guided objectives, Humansplat achieves more detailed and higher fidelity results against GTA GTA:2024 and LGM LGM:2024 , as illustrated in Fig. 4. Results on in-the-wild images demonstrate our strong generalization capability. Although TeCH TECH:2023 also generates high quality image, its SDS optimization often results in overly saturated multi-face outcomes poole2022dreamfusion ; wang2023score . Moreover, while TeCH delivers highest PSNR on the Thuman2.0 dataset, the LPIPS metric in Tab. 1, confirms that our method delivers more realistic results. Notably, our model provides superior textures on both the front and invisible regions, highlighting the effectiveness of incorporating appearance and geometric priors.
4.3 Ablation Study
Latent Reconstruction Transformer. HumanSplat leverages the Stable Diffusion latent space for feature extraction. Experimental results demonstrate that utilizing the compact latent space for reconstructions leads to better rendering quality compared to using the pixel space, as detailed in Tab. 2(c) (a). We also find that patching in the time dimension for latent decoding by temporal VAE SVD:2023 adversely affects the quality, so we employ the vanilla VQ-VAE VQGAN:2023 decoder in this work.
[…]
Method Category THuman2.0 THuman2.0:2021 Twindom Twindom Rec. Time
[…]
PIFu PIFu:2019 18.093 0.911 0.137 - - - $\sim$30 Sec
LGM LGM:2024 20.013 0.893 0.116 19.840 0.851 0.292 $\sim$9.5 Sec
GTA GTA:2024 18.050 - - 17.669 0.741 0.418 $\sim$5.6 Min
SIFU SIFU:2024 22.025 0.921 0.084 19.714 0.832 0.312 $\sim$5.7 Min
SIFU SIFU:2024 22.102 0.923 0.079 - - - $\sim$5.7 Min
Magic123 magicanimate:2023 14.501 0.874 0.145 - - - $\sim$1 Hours
HumanSGD HumanSGD:2023 17.365 0.895 0.130 - - - $\sim$7 Min
TeCH TECH:2023 25.211 0.936 0.083 21.192 0.884 0.188 $\sim$4.5 Hours
HumanSplat (Ours) 24.033 0.918 0.055 23.346 0.913 0.125 $\sim$9.3 Sec
Figure 4: Qualitative comparison of ours against TeCH TECH:2023 , GTA GTA:2024 and LGM LGM:2024 on Thuman2.0 THuman2.0:2021 , Twindom Twindom and in-the-wild images. Our method achieves the highest quality. Note that TeCH achieves clearer results but fails to preserve the face ID.
Human Geometric Prior. To evaluate the robustness of HumanSplat to SMPL parameters, we employed various initialization strategies for the geometry. The results of these different methods are presented in Table 2(c) (b). In cases where the SMPL initialization involved extreme randomness, our model demonstrated robustness against SMPL estimation errors, as our approach of treating human priors as keys and values. Furthermore, the projection window $k_{win}$ ensures that parameters estimated from a single view detector PIXIE:2021 , as well as results from the multi-view version of SMPLify Pavlakos2019_smplifyx , remain within acceptable tolerances. The effectiveness of the human geometric prior is also demonstrated qualitatively in Figure 5 (a).
Window Size. As analyzed in Sec. 3.4.2, human priors are crucial for the robustness of the human reconstruction model. As demonstrated in Tab. 2(c) (c), the PSNR metric is significantly improved from 22.625db to 24.293db. When the window size $K_{win}$ is set to 1 (Project Only), the model struggles to handle loose clothing. We also explore how the performance of the model varies with different window sizes by setting $K_{win}=2$, in a latent space with a resolution of $64\times 64$.
Semantics-guided Training Objectives. Thanks to the incorporation of semantic-guided hierarchical loss, HumanSplat can generate textures of higher quality for both the face and hand, while also promote the convergence of the training process, as depicted in Figure 5 (a) and (b). Furthermore, Figure 5 (c) demonstrates that HumanSplat further produces 3D consistent segmentation results, which have potential applications in areas such as editing and specific generation of individual parts of 3D humans.
Figure 5: (a) Qualitative evaluation of human prior and hierarchical loss $\mathcal{L_{H}}$. PSNR/SSIM values are shown at the top of each image, presenting the effectiveness of the two strategies in the full pipeline. (b) LPIPS scores during training on the 2K2K validation set. (c) HumanSplat can provide 3D consistent segmentation results for potential applications.
Table 2: Ablation study of several proposed designs on 2K2K  2k2k dataset.
(a) Reconstruction Space.
[…]
Pixel Space 23.125 0.892 0.066
Temporal VAE SVD:2023 24.133 0.902 0.047
Ours rombach2022high 24.293 0.915 0.040
(b) Geometric Prior Initialization.
[…]
Random 19.894 0.876 0.366
Muti-view 24.633 0.935 0.025
Ours PIXIE:2021 24.293 0.915 0.040
(c) Different project methods.
[…]
w/o Human Prior 22.635 0.893 0.182
Project Only 23.333 0.916 0.057
Ours($K_{win}=2$) 24.293 0.915 0.040
5 Conclusion
We present HumanSplat, a pioneering generalizable Human Gaussian Splatting network that derives 3D Gaussian Splatting properties from a single image. This model integrates generative diffusion and latent reconstruction transformer models with human structure priors, enhanced by a tailored semantic-aware hierarchical loss. These innovations achieve realistic and detailed reconstructions in a fast feed-forward manner without any optimization or fine-tuning, especially in the important focal areas such as the face and hands. Extensive experiments demonstrate that HumanSplat surpasses existing state-of-the-art methods in both quality and efficiency, including robust performance in handling challenging poses and loose clothing. This capability opens up a broad spectrum of potential applications.
Limitations and Future Works. Our methods have some limitations that can be addressed in future works: (a). While effective with most attire, the model may struggle with intricate garments and accessories. Future enhancements should incorporate 2D data and innovative training strategies, which is a promising direction for improving the handling of diverse and unusual clothing styles. (b). Although already efficient, increasing the computational speed by constraining Gaussian’s number or combining compact 3DGS with texture representation could further facilitate real-time applications, particularly on mobile devices. (c). Animating the reconstructed human models currently requires post-processing. Future work could introduce canonical space to reconstruct animatable avatars directly, streamlining the process and improving efficiency.
Appendix A Implementation Details
A.1 Additional ovel View Synthesizer Details
The conditions of Difffusion model. SVD and SV3D SVD:2023 ; voleti2024sv3d replace CLIP text embeddings CLIP:2021 with the image embedding of the conditioning. Compared to text prompts, image prompts provide information more accurately and directly, and also save time that would otherwise be spent on text inversion or image caption  TECH:2023 ; SIFU:2024 .
Transfomer Block. The CLIP embedding matrix from the conditioning image is fed into the cross-attention layers of each transformer block, acting as the key and value, while the feature at that layer serves as the query. Moreover, the camera trajectory and the diffusion noise timestep are incorporated into the residual blocks.
[…]
A.2 Additional Latent Reconstruction Transformer Details
The network architecture of the proposed latent reconstruction transformer is shown in Fig. 6. The framework is composed of a latent encoder with intra- and inter-attention and a Gaussian parameter prediction module. It takes initial multi-view latent embeddings as input and performs projection-aware attention to aggregate human geometric information. From each output token, we decode the attributes of pixel-aligned Gaussians in the corresponding patch with a Conv $1\times 1$ layer.
Figure 6: Detailed network architecture of latent reconstruction transformer.
A.3 Additional Training Details
[…]
Table 3: Evaluation of different window sizes $K_{win}$ on the 2K2K evaluation dataset.
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
w/o Human Prior 22.635 (-1.658) 0.893 (-0.022) 0.182 (+0.142)
$K_{win}=1$ 23.333 (-0.960) 0.916 (-0.001) 0.057 (+0.017)
$K_{win}=3$ 24.383 (+0.089) 0.931 (+0.016) 0.041 (+0.001)
$K_{win}=4$ 24.013 (-0.280) 0.922 (+0.007) 0.048 (+0.008)
$K_{win}=2$ (Ours) 24.293 0.915 0.040
Figure 7: Qualitative 3D Gaussian Splatting results of diversified evaluation dataset. (a) Input Image. (b) Novel view Rendering Results.
Appendix B More Results.
In Tab. 3, we present results for the 2K2K dataset across various window sizes. The results show that employing a window size of $k_{\text{win}}=1$ significantly enhances performance, improving the PSNR metric from 22.635 db to 23.333 db. Further refinement with a window size of $k_{\text{win}}=3$ continues to enhance the results, whereas using $k_{\text{win}}=4$ does not yield additional improvements. After considering both computational complexity and metrics, we selected a window size of $K_{win}=2$.
As for novel view and novel pose rendering, more results are shown in Fig. 7 and Fig. 8. For a single input image, we use AnimateAnyone AnimateAnyone:2023 for novel pose synthesis and HumanSplat for novel view synthesis. This functionality enables dynamic timestamps and real-time rendering of novel views, thereby enhancing immersive virtual exploration. Please also kindly refer to our supplementary video for more results.
Appendix C Broader Impacts
While our model demonstrates the capability to reconstruct photo-realistic 3D avatars, it also introduces potential risks such as privacy violations. To mitigate these concerns, it is imperative to establish robust ethical guidelines and legal frameworks. This necessitates a collaborative effort among researchers, developers, and policymakers to promote the responsible use of this technology and safeguard against its potential misuse.
Figure 8: Qualitative 4D Gaussian Splatting results on In-the-wild images, including novel view and pose rendering results.